How many moles of hydrogen fluoride will be produced if 243 grams of ammonia(NH3) reacts with fluorine?
1
Expert's answer
2019-06-03T05:49:49-0400

The Stoichiometric equation for the reaction is

2NH 3 + 6F 2 → 6HF + 2NF 3

Thus according to above equation 2 moles of NH 3 reacts with 6 moles of fluorine to form 6 moles

of HF

Moles of NH 3 = (243/17) = 14.29 moles

2 moles of NH 3 = 6 moles of HF

1 mole of NH 3 = 3 moles of HF

14.29 moles of NH 3 = 3*14.29 mole of HF

Thus 42.87 moles of HF are produced.